# Modeling of Memory Mechanisms in Cerebral Cortex and Simulation of Storage Performance

### v1

@@ 1 Introduction
The brain, as a complex information processing system, remains incompletely understood, particularly its memory mechanisms. Exploring how the brain encodes, stores, and retrieves information has always been a significant challenge in memory research. Modern neuroimaging technology is rapidly advancing, and through methods such as fluorescent tagging and two-photon imaging, humans are gradually unveiling the neural system’s connectivity and other details. However, such analyses can only offer limited insights, and we are still unfamiliar with the neural mechanisms and information flow directions underlying memory activities. From a neurobiological perspective, memory refers to neural system activities or physical changes in neuronal connections triggered by external stimuli or brain states chaudhuri_computational_2016 . Neurobiological research on memory mechanisms focuses on synaptic plasticity martin_synaptic_2000 ; abraham_is_2019 ; jeong_synaptic_2021 ; humeau_next_2019 , neurotransmitters de_rossi_neuronal_2020 ; feld_neurochemical_2020 , and electrochemical signals wang_single-vesicle_2022 ; otero_exploring_2022 at the cellular and molecular levels, while cognitive psychology likens memory to an information processing system responsible for encoding, storing, and retrieving information. These theories provide a reasonable starting point to map the brain’s coarse-scale organization using functional imaging technologies including EEG and fMRI kriegeskorte_cognitive_2018 .
@@
Minimum Functional Unit
Data are stored in binary form, with hard drives using polarity of magnetic particles, charges in capacitors, etc., to represent 1s or 0s.
In neural cells, membrane possesses resting and action potentials. In resting state, signal is inhibited; if activated, it generates an action potential, conducting signal.
Storage Device
An HDD records data by changing the distribution of the two polarities of magnetic particles using an electromagnet on read/write head; an SDD stores 0s and 1s by altering the number of electrons in the floating gate layer through the application of an electric field.
Memory refers to neural system activities or physical changes in neuronal connections triggered by external stimuli or brain states chaudhuri_computational_2016 . Many studies have begun to define engrams as basic units of memory, but there are unresolved issues such as how structure of engrams affects quality of memory, how multiple engrams interact with each other, and how engrams change over time josselyn_memory_2020 .
Data Allocation
Operating systems allocate disk space for files with blocks as the basic unit, using methods such as contiguous allocation, linked allocation, and indexed allocation.
We have a very limited understanding of how a single memory item is structurally stored at the level of biological neurons, and how multiple memory items are allocated among neural cells (groups) in cerebral cortex.
Data Access
Control circuits translate logical addresses into physical addresses. Read/write head of an HDD moves to the cylinder, track, and sector where the data are located for data access; for an SDD, controller directly performs read and write operations on storage units at physical address.
When brain stores or recalls a memory, we know little about how it locates engram of memory content among vast number of neural cells in cerebral cortex, replacement of old memories with new ones, and neuronal network mechanisms during recall.
@@
The brain can be conveniently represented as a network of neurons and their interconnections, making graph theory a mathematical tool for studying its structure and functional systems. We utilize directed graphs from graph theory to model and simulate the algorithmic layer of the brain’s memory system, aiming to align anatomical and physiological data. As a branch of mathematics, graph theory provides tools for processing and analyzing network structures. Tracing back to Euler’s solution to the Seven Bridges of Konigsberg problem in the 18th century gao_graph_2009 , it has been applied such as to design solutions for the Traveling Salesman Problem(TSP) cook_traveling_2007 , construct knowledge graphs ji_survey_2021 , and create databases using graph structures jouili_empirical_2013 , where graphs usually serve as structured representations of data or knowledge, inherently lacking dynamic behavior, with their functionality reliant on externally applied algorithms.
@@ 2.1 Classical Graph Theory
The TSP is a classic NP-hard problem, which involves finding the shortest path that visits each node exactly once and returns to the starting point in a complete graph cook_traveling_2007 . It has a wide impact in areas including logistics, flight route planning, network design, and bioinformatics. Due to its high computational complexity, various methods have been proposed to solve or approximate it, such as branch and bound, nearest neighbor, simulated annealing, and genetic algorithms. These methods essentially act on a completely known graph, possessing a “God’s eye” view.
One of the earliest large-scale applications of classical graph theory in computer science is graph databases. Representing modern applications such as social networks in graph form is quite natural jouili_empirical_2013 . Graph databases represent data or patterns as graphs, with data structures modeled as directed, possibly labeled graphs. Data operations are represented by graphs, and operation-oriented and type constructors can be defined on graph structures along with appropriate integrity constraints angles_survey_2008 . As defined, the implementation of graph databases relies on external algorithms for different operations, such as graph construction, exhibiting a “God’s eye” view.
Another important application of classical graph theory is the knowledge graph, i.e., a structured representation of facts, entities, relationships, and semantic descriptions. Knowledge is interconnected, with knowledge nodes forming a graph network to integrate various knowledge domains ji_survey_2021 . Knowledge graphs possess a complete topology and leverage their structure to integrate knowledge, providing a “God’s eye” view. However, the graph remains static after information updates.
@@ 2.2 Graph Theory and Deep Learning
Graphs are now widely used in deep learning. Tasks involving models on graphs can categorized into two types: Node-centered tasks are associated with individual nodes in a graph and include node classification, link prediction, and node recommendation. Graph-centered tasks are related to the entire graph and include graph classification, estimation of various graph properties, and graph generation zhang_deep_2020 . A well-known application, the Graph Convolutional Network (GCN), inputs graphs to convolutional networks for training jia_multimodal_2023 . Operations in deep learning involve manipulating graph data, relying on a “God’s eye” view, while the graph remains static.
Graph theory is extensively used in deep learning applications. In computer vision, node encoding methods include both non-neural networks, such as traditional graph embedding and probabilistic graph models, and neural networks, including Graph Recurrent Neural Networks, GCNs, and variants of Graph Neural Networks jiao_graph_2022 . Again, these methods process graph data, and the graph itself is static.
@@ 2.3 Multi-agent Systems
Multi-agent systems have gained widespread attention in recent years. These decentralized systems accomplish a task through the collaborative effort of many small entities, each with independent behavior. The small entities make distributed independent decisions with individual tasks assigned to autonomous entities called agents, each determining the correct action for solving the task based on their inputs. Individual agents only have partial information and communicate with their neighbors. These are applied in areas such as computer networks, drones, and robot swarms dorri_multi-agent_2018 .
Multi-agent systems are inspired by the natural world, such as in ant colonies and bee societies leitao_bio-inspired_2012 . Each ant in a colony achieves complex tasks such as finding food, building nests, and collective defense through simple pheromone communication and collaborative behavior. Each bee communicates effective foraging information, temperature regulation, and new nest site selection through precise dances and collaboration. Through local interactions and simple rules, they achieve complex global behavior, providing inspiration and reference for computer science and promoting the development and innovation of artificial multi-agent systems.
Multi-agent systems are extensively used in robotic systems and to some extent can solve issues such as local positioning, obstacle detection, path planning, and navigation in multi-robot systems rasheed_review_2022 . For example, a collision avoidance problem was solved for multiple robots in a decentralized, distributed framework, where input data were only collected from onboard sensors long_towards_2018 . Each robot is an agent that shares its decision with all other robots, thereby achieving an optimal strategy among them.
@@ 3 Modeling Based on Directed Graphs
Many researchers have modeled memory. The Hopfield network hopfield_neural_1982 is a type of recurrent neural network that can be abstracted as a directed complete graph, with many variants krotov_dense_2016 ; kobayashi_chaotic_2017 ; oku_pseudo-orthogonalization_2013 ; mungai_chunking_2017 ; mungai_semantic_2017 ; shriwas_multi-modal_2019 . Bidirectional associative memory networks kosko_bidirectional_1988 implement bidirectional associative functionality and have undergone many modifications zhao_synchronization_2021 ; kosko_bidirectional_2021 ; singh_multilayer_2017 ; cholet_bidirectional_2019 . Structural learning has also been widely used gershman_computational_2017 ; wei_hierarchical_2007 ; kimoto_mixed_2004 ; dekhtyarenko_systematic_2005 ; he_constructing_2019 , with some models considering the impact of memory retrieval on memory storage gershman_computational_2017 . There are also sparse associative memory models hoffmann_sparse_2019 and models based on dictionary learning mazumdar_associative_2017 ; tosic_dictionary_2011 . Models based on graph theory, such as clique-based models gripon_sparse_2011 and competition graph models jiang_storing_2015 ; mofrad_neural_2021 , use graph theory and network flow theory to simulate human memory. These models have made progress in simulating human memory, but nevertheless have limitations. For instance, dense associative memorykrotov_dense_2016 lacks sufficient simulation of neuronal structure and function; the Hopfield hopfield_neural_1982 and BAM networks kosko_bidirectional_1988 have a gap in fully connected networks and biological realism; and clique-based models gripon_sparse_2011 process memory in a way not entirely consistent with actual mechanisms in the human brain. Many models face challenges in hardware implementation, computational resource requirements, and data processing efficiency.
Electronic components such as memristors chua_memristor-missing_1971 ; tour_fourth_2008 ; itoh_memristor_2008 have also been proposed to model memory. Their resistance changes when electric current passes through, but remains constant if current flow stops. This memory of resistance makes them like biological neural synapses, and they are often used to construct memory computing models sun_memristor-based_2019 ; sun_quasi-ideal_2014 ; an_realizing_2019 ; li_novel_2019 ; chen_adaptive_2018 . However, memristor-based memory models have shortcomings, such as only simulating a few neuronal synapses, and thus differing from biological reality sun_memristor-based_2019 , or focusing on circuit analysis and neglecting biological emulation chen_adaptive_2018 .
@@ 6.1 Generation of Directed Graph
                   for neighbor normal-$\leftarrow$ neighbors[0] to neighbors[neiLen - 1] do
@@ 6.5 Capacity Testing
Size of Directed Graph Network
Number of Connections Between Nodes in Connected Subgraph of Instance
Average Number of Nodes Constituting Path
30 × 10
364.459
54.439
30 × 20
813.775
83.721
30 × 30
1385.789
113.686
@@ 6.6 Capacity Comparison
We compare the capacity of the proposed model with that of the Hopfield network measured in experiments. Like the criterion of $\alpha\geq 90\%$ set in this paper, in the Hopfield network, a retrieved vector is considered successfully remembered if its similarity with the original vector exceeds 90%. The memory efficiency of the Hopfield network is optimal when each binary bit of a storage instance is independently generated with a 50% probability of being one of two values mceliece_capacity_1987 , which corresponds to randomly activating 50% of the edge nodes in the proposed model. To control variables, other parameter settings remained the same as in Section 6.5, but we randomly activated 50% of the edge nodes. At the same time, $\rho$ for the probe vector used in the Hopfield network was the same as $\rho_{max}$.
@@ 6.7 Capacity Testing of Networks with Six Layers of Depth
Network Width
Number of Connections Between Nodes in Connected Subgraph of Instance
Average Number of Nodes Constituting Path
@@ 6.8 Chain Awakening Experiment with Long Chains
Many brain activities, such as sequential working memory, are related to temporal information xie_geometry_2022 . A long-chain sequential awakening experiment examined a series of instances with sequential associations, where the output component of a memory instance is the input for the following instance, i.e., it acts as an awakening cue. The experiment observed whether such a chain-style awakening can be successful. Multiple vectors of the same length were randomly generated and combined in pairs to form memory instances, with the output vector (D component) of one instance serving as the input vector (A component) of the next. The method of uploading memory instances is similar to Figure 19 and illustrated in Figure 21, the edge nodes used by each instance do not overlap, indicating that the instances do not share input or output nodes.

### v2

@@ 1 Introduction
The brain, as a complex information processing system, remains incompletely understood, particularly its memory mechanisms. Exploring how the brain encodes, stores, and retrieves information has always been a significant challenge in memory research. Modern neuroimaging technology is rapidly advancing, and through methods such as fluorescent tagging and two-photon imaging, humans are gradually unveiling the neural system’s connectivity and other details. However, such analyses can only offer limited insights, and we are still unfamiliar with the neural mechanisms and information flow directions underlying memory activities. From a neurobiological perspective, memory refers to neural system activities or physical changes in neuronal connections triggered by external stimuli or brain states [1]. Neurobiological research on memory mechanisms focuses on synaptic plasticity [2, 3, 4, 5], neurotransmitters [6, 7], and electrochemical signals [8, 9] at the cellular and molecular levels, while cognitive psychology likens memory to an information processing system responsible for encoding, storing, and retrieving information. These theories provide a reasonable starting point to map the brain’s coarse-scale organization using functional imaging technologies including EEG and fMRI [10].
@@
Minimum Functional Unit Data are stored in binary form, with hard drives using polarity of magnetic particles, charges in capacitors, etc., to represent 1s or 0s. In neural cells, membrane possesses resting and action potentials. In resting state, signal is inhibited; if activated, it generates an action potential, conducting signal.
Storage Device An HDD records data by changing the distribution of the two polarities of magnetic particles using an electromagnet on read/write head; an SDD stores 0s and 1s by altering the number of electrons in the floating gate layer through the application of an electric field. Memory refers to neural system activities or physical changes in neuronal connections triggered by external stimuli or brain states [1]. Many studies have begun to define engrams as basic units of memory, but there are unresolved issues such as how structure of engrams affects quality of memory, how multiple engrams interact with each other, and how engrams change over time [11].
Data Allocation Operating systems allocate disk space for files with blocks as the basic unit, using methods such as contiguous allocation, linked allocation, and indexed allocation. We have a very limited understanding of how a single memory item is structurally stored at the level of biological neurons, and how multiple memory items are allocated among neural cells (groups) in cerebral cortex.
Data Access Control circuits translate logical addresses into physical addresses. Read/write head of an HDD moves to the cylinder, track, and sector where the data are located for data access; for an SDD, controller directly performs read and write operations on storage units at physical address. When brain stores or recalls a memory, we know little about how it locates engram of memory content among vast number of neural cells in cerebral cortex, replacement of old memories with new ones, and neuronal network mechanisms during recall.
@@
The brain can be conveniently represented as a network of neurons and their interconnections, making graph theory a mathematical tool for studying its structure and functional systems. We utilize directed graphs from graph theory to model and simulate the algorithmic layer of the brain’s memory system, aiming to align anatomical and physiological data. As a branch of mathematics, graph theory provides tools for processing and analyzing network structures. Tracing back to Euler’s solution to the Seven Bridges of Konigsberg problem in the 18th century [12], it has been applied such as to design solutions for the Traveling Salesman Problem(TSP) [13], construct knowledge graphs [14], and create databases using graph structures [15], where graphs usually serve as structured representations of data or knowledge, inherently lacking dynamic behavior, with their functionality reliant on externally applied algorithms.
@@ 2.1 Classical Graph Theory
The TSP is a classic NP-hard problem, which involves finding the shortest path that visits each node exactly once and returns to the starting point in a complete graph [13]. It has a wide impact in areas including logistics, flight route planning, network design, and bioinformatics. Due to its high computational complexity, various methods have been proposed to solve or approximate it, such as branch and bound, nearest neighbor, simulated annealing, and genetic algorithms. These methods essentially act on a completely known graph, possessing a “God’s eye” view.
One of the earliest large-scale applications of classical graph theory in computer science is graph databases. Representing modern applications such as social networks in graph form is quite natural [15]. Graph databases represent data or patterns as graphs, with data structures modeled as directed, possibly labeled graphs. Data operations are represented by graphs, and operation-oriented and type constructors can be defined on graph structures along with appropriate integrity constraints [16]. As defined, the implementation of graph databases relies on external algorithms for different operations, such as graph construction, exhibiting a “God’s eye” view.
Another important application of classical graph theory is the knowledge graph, i.e., a structured representation of facts, entities, relationships, and semantic descriptions. Knowledge is interconnected, with knowledge nodes forming a graph network to integrate various knowledge domains [14]. Knowledge graphs possess a complete topology and leverage their structure to integrate knowledge, providing a “God’s eye” view. However, the graph remains static after information updates.
@@ 2.2 Graph Theory and Deep Learning
Graphs are now widely used in deep learning. Tasks involving models on graphs can categorized into two types: Node-centered tasks are associated with individual nodes in a graph and include node classification, link prediction, and node recommendation. Graph-centered tasks are related to the entire graph and include graph classification, estimation of various graph properties, and graph generation [17]. A well-known application, the Graph Convolutional Network (GCN), inputs graphs to convolutional networks for training [18]. Operations in deep learning involve manipulating graph data, relying on a “God’s eye” view, while the graph remains static.
Graph theory is extensively used in deep learning applications. In computer vision, node encoding methods include both non-neural networks, such as traditional graph embedding and probabilistic graph models, and neural networks, including Graph Recurrent Neural Networks, GCNs, and variants of Graph Neural Networks [19]. Again, these methods process graph data, and the graph itself is static.
@@ 2.3 Multi-agent Systems
Multi-agent systems have gained widespread attention in recent years. These decentralized systems accomplish a task through the collaborative effort of many small entities, each with independent behavior. The small entities make distributed independent decisions with individual tasks assigned to autonomous entities called agents, each determining the correct action for solving the task based on their inputs. Individual agents only have partial information and communicate with their neighbors. These are applied in areas such as computer networks, drones, and robot swarms [20].
Multi-agent systems are inspired by the natural world, such as in ant colonies and bee societies [21]. Each ant in a colony achieves complex tasks such as finding food, building nests, and collective defense through simple pheromone communication and collaborative behavior. Each bee communicates effective foraging information, temperature regulation, and new nest site selection through precise dances and collaboration. Through local interactions and simple rules, they achieve complex global behavior, providing inspiration and reference for computer science and promoting the development and innovation of artificial multi-agent systems.
Multi-agent systems are extensively used in robotic systems and to some extent can solve issues such as local positioning, obstacle detection, path planning, and navigation in multi-robot systems [22]. For example, a collision avoidance problem was solved for multiple robots in a decentralized, distributed framework, where input data were only collected from onboard sensors [23]. Each robot is an agent that shares its decision with all other robots, thereby achieving an optimal strategy among them.
@@ 3 Modeling Based on Directed Graphs
Many researchers have modeled memory. The Hopfield network [24] is a type of recurrent neural network that can be abstracted as a directed complete graph, with many variants [25, 26, 27, 28, 29, 30]. Bidirectional associative memory networks [31] implement bidirectional associative functionality and have undergone many modifications [32, 33, 34, 35]. Structural learning has also been widely used [36, 37, 38, 39, 40], with some models considering the impact of memory retrieval on memory storage [36]. There are also sparse associative memory models [41] and models based on dictionary learning [42, 43]. Models based on graph theory, such as clique-based models [44] and competition graph models [45, 46], use graph theory and network flow theory to simulate human memory. These models have made progress in simulating human memory, but nevertheless have limitations. For instance, dense associative memory[25] lacks sufficient simulation of neuronal structure and function; the Hopfield [24] and BAM networks [31] have a gap in fully connected networks and biological realism; and clique-based models [44] process memory in a way not entirely consistent with actual mechanisms in the human brain. Many models face challenges in hardware implementation, computational resource requirements, and data processing efficiency.
Electronic components such as memristors [47, 48, 49] have also been proposed to model memory. Their resistance changes when electric current passes through, but remains constant if current flow stops. This memory of resistance makes them like biological neural synapses, and they are often used to construct memory computing models [50, 51, 52, 53, 54]. However, memristor-based memory models have shortcomings, such as only simulating a few neuronal synapses, and thus differing from biological reality [50], or focusing on circuit analysis and neglecting biological emulation [54].
@@ 6.1 Generation of Directed Graph
                   for neighbor $\leftarrow$ neighbors[0] to neighbors[neiLen - 1] do
@@ 6.5 Capacity Testing
Size of Directed Graph Network Number of Connections Between Nodes in Connected Subgraph of Instance Average Number of Nodes Constituting Path
30 × 10 364.459 54.439
30 × 20 813.775 83.721
30 × 30 1385.789 113.686
@@ 6.6 Capacity Comparison
We compare the capacity of the proposed model with that of the Hopfield network measured in experiments. Like the criterion of $\alpha\geq 90\%$ set in this paper, in the Hopfield network, a retrieved vector is considered successfully remembered if its similarity with the original vector exceeds 90%. The memory efficiency of the Hopfield network is optimal when each binary bit of a storage instance is independently generated with a 50% probability of being one of two values [55], which corresponds to randomly activating 50% of the edge nodes in the proposed model. To control variables, other parameter settings remained the same as in Section 6.5, but we randomly activated 50% of the edge nodes. At the same time, $\rho$ for the probe vector used in the Hopfield network was the same as $\rho_{max}$.
@@ 6.7 Capacity Testing of Networks with Six Layers of Depth
Network Width Number of Connections Between Nodes in Connected Subgraph of Instance Average Number of Nodes Constituting Path
@@ 6.8 Chain Awakening Experiment with Long Chains
Many brain activities, such as sequential working memory, are related to temporal information [56]. A long-chain sequential awakening experiment examined a series of instances with sequential associations, where the output component of a memory instance is the input for the following instance, i.e., it acts as an awakening cue. The experiment observed whether such a chain-style awakening can be successful. Multiple vectors of the same length were randomly generated and combined in pairs to form memory instances, with the output vector (D component) of one instance serving as the input vector (A component) of the next. The method of uploading memory instances is similar to Figure 19 and illustrated in Figure 21, the edge nodes used by each instance do not overlap, indicating that the instances do not share input or output nodes.